# Dynamic Fog Computing for Enhanced LLM Execution in Medical Applications

### v1

@@ 6 Conclusion
SpeziLLM’s primary contribution is its uniform, LLM-agnostic interface, which facilitates transparent transitions between inference environments. This approach offers a viable solution to the challenges of LLM execution in sensitive and resource-constrained environments [21, 22, 23], facilitating the segmentation of extensive LLM tasks into smaller components and the distribution of such segments across complexity- and trust-matched layers. Lightweight models, such as Llama 2, can efficiently handle simpler tasks within the local and fog layers, while more complex tasks can be processed in the cloud. Future developments in efficient LLM execution with limited memory will further refine SpeziLLM’s performance and applicability.
@@
The research was partially funded by a German Academic Exchange Service (DAAD) ”Internationale Forschungsaufenthalte für Informatikerinnen & Informatiker” scholarship (91886835) and the Bavaria California Technology Center (BaCaTec) project fund (Nr. 10 2023-2). We thank the Stanford Byers Center for Biodesign for supporting this project and our digital health research.

### v2

@@ 6 Conclusion
SpeziLLM’s primary contribution is its uniform, LLM-agnostic interface, which facilitates transparent transitions between inference environments. This approach offers a viable solution to the challenges of LLM execution in sensitive and resource-constrained environments [21, 22, 23, 24], facilitating the segmentation of extensive LLM tasks into smaller components and the distribution of such segments across complexity- and trust-matched layers. Lightweight models, such as Llama 2, can efficiently handle simpler tasks within the local and fog layers, while more complex tasks can be processed in the cloud. Future developments in efficient LLM execution with limited memory will further refine SpeziLLM’s performance and applicability.
@@
The research was partially funded by a German Academic Exchange Service (DAAD) ”Internationale Forschungsaufenthalte für Informatikerinnen & Informatiker” scholarship (91886835) and the Bavaria California Technology Center (BaCaTec) project fund (Nr. 10 2023-2). We thank the Stanford Mussalem Center for Biodesign for supporting this project and our digital health research.